OpenStudy (kpop4life123):

HELP IM STUCK IN THIS PROBLEM >_< I WILL FAN AND MEDAL!!!!! Find the value of the following expression: (2^8 ⋅ 3^–5 ⋅ 6^0)^–2 ⋅ (3^-2 / 2^3)^4 ⋅ 2^28 Write your answer in simplified form. Show all of your steps.

OpenStudy (kpop4life123):

@bbypanda16

OpenStudy (kpop4life123):

@adll23

OpenStudy (michele_laino):

step by step, we have: \[\huge {\left( {{2^3}} \right)^4} = {2^{3 \cdot 4}} = ...?\]

OpenStudy (kpop4life123):

@Michele_Laino 4096?

OpenStudy (michele_laino):

correct! better is \( 2^{12}\)

OpenStudy (michele_laino):

now,, what is: \[\huge {\left( {{3^{ - 2}}} \right)^4} = {3^{\left( { - 2 \cdot 4} \right)}} = ...?\] please give your answer in form of power

OpenStudy (bbypanda16):